2025 TLC

  • Reset south facing laser and see why it’s hitting 10cm
  • 2x Cr2032 batteries for bios
  • More pipe locations for bamboo summer mounts
  • Dozier plot sign
  • Fix cimel bolts
  • South laser has diurnal cycle. Could it be caused by thermal expansion and torsion of the railing? Test: identify the laser’s target location, flatten it out and observe if noise persists.
  • Clean and refill drinking water
  • Add Dozier Site to google maps
    • Awaiting response from Google
  • Add site to Open Street Map
  • Remove DRI meltwater sampler
  • Shopping
    • Paper towels (careful. The mouse like this as bedding. Need to have it in a mouse-safe box)
    • Bleach
    • Running low on some of the shrink wrap sizes. Get more. Maybe colored ones
    • Vacuum: Get a vacuum filter (shop-vac 6 gallon “quiet series”
    • Get connectors (don’t even have to be waterproof) to easily plug things in and out example
    • Better ethernet crimping tool. And a tester!
    • Soldering Iron: It seems like the soldering iron cable might have a loose contact. Voltage from station is good. Either
      – redo the cable or
      – replace the “pencil” (iron/heater) (TC201T). Just an example: [amazon link](https://www.amazon.com/Weller-TC201T-Solder-Pencil-Wtcpt/dp/B000ICCZ4O)
    • Get hand disinfectant for mouse traps
    • Will run low on solder soon
    • Mouse mitigation. See below
  • Tear down infrasound
  • Take detailed pictures of instrument
  • Fix lymimister joint (leaking at pan)
  • Find better place for JPL Radar
  • Move the spare eppleys to snarl
  • Update logger 2 program:
    • remove snow temperatures from logger 2
    • The new buckets are 100 ml, not 80 ml tips. This is not reflected in the (software) calibration → need up update logger2 program and correct past measurements
  • Need to correct past lysimeter measurements in database since buckets were swapped
  • Need to add README about correcting past measurements in raw files since buckets were swapped
  • Fix issue with missing tipping bucket measurements followed by doubling of measurements in next scan
    • Solution: change delay command to > 5 seconds in logger program
  • Verify physical bucket calibration and reed switch/magnet clearance
    • Figure out what leads to tip counts that are integer multiples of 256
  • At least one lightbulb needs replacing. Probably replace all with LEDs to reduce fire risk etc
  • Attempt fixing camera
  • Test all the ethernet connectors to local network
  • Rodent/mouse mitigation
    • Purchase + install new/second ultrasonic emitter
    • Remove or seal everything that could be used as bedding, e.g. paper towels at end of TLC(tote box?)
    • Absolutely no eating in the bunker during TLC
    • Remove all bedding material such as styrofoam cardboard, packing material
    • Consider/find new/additional mitigation options.
      • Get a cat and/or a hawk and/or snakes (I hope there is no way of tracing back who adds items to this list)
      • Is there a chance to prevent them from entering / sealing the door? I saw one running out through the cable conduits.
        • Seems like we could wire-mesh or steel-wool around the cables.
      • (Better) bait for the traps?
      • Thoroughly clean/vacuum at the end of TLC
  • logger 3 os updated
  • Steel wool (mouse)
  • Totebox (mouse protection)
  • fix tour on webcam
  • Clean bunker computer keyboards